OpenStudy (anonymous):

if dy/dx is inversely proportional to yand y=2 when x=0 and y=4 when x=2, find y when x=3. does the mod in ln|x| go away??

OpenStudy (anonymous):

given dy/dx = k/y where k is constant of proportionality

OpenStudy (anonymous):

so separating the variables we have ydy=kdx

OpenStudy (anonymous):

yeah and then i got x=kln|y|+c

OpenStudy (anonymous):

no integrating ydy=kdx (y^2)/ 2= kx +c where c is integrating constant

OpenStudy (anonymous):

whenx x=0 y=2 implies c=2

OpenStudy (anonymous):

okay conitnue?

OpenStudy (anonymous):

now when x=2 then y=4 that means 16/2 =2k+2 implies k=3

OpenStudy (anonymous):

OH WAIT ONE SECOND. WROTE QUESTOIN WRONG. SORRY

OpenStudy (anonymous):

its dx/dy=k/y not dy/dx

OpenStudy (anonymous):

hence (y^2 )/2 = 3x +2 or y^2=6x+4

OpenStudy (anonymous):

but you have posted dy/dx

OpenStudy (anonymous):

yeah my bad .. sorry i typed wrong its dx/dy

OpenStudy (anonymous):

then dx/dy = k/y is it correct

OpenStudy (anonymous):

again sepearte the variables dx = kdy/y

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

x=ln|y|k +c ?

OpenStudy (anonymous):

or is it just lny is there mod ?

OpenStudy (anonymous):

x= k(ln|y|-ln|c|) instead of C we can use ln|c|

OpenStudy (anonymous):

howcome??

OpenStudy (anonymous):

ln|y/c|=x/k

OpenStudy (anonymous):

for constants we can use c or logc or sinc as per depending on the question to make work easier

OpenStudy (anonymous):

or y=ce^(x/k) when x= 0 y=2 implies c=2 hence y=2* e^(x/k)

OpenStudy (anonymous):

ln|y/c|=x/k or k = x/ ln|y/c| now when x=2 y=4 and alsoc c=2 hence k= 2/ ln|4/2|= 2/ln2